# Towards Generalized Video Quality Assessment: A Weak-to-Strong Learning Paradigm

### v3

@@ 4.4 Experimental Results
Iterative W2S Training. We observe consistent improvements across both in-domain and OOD datasets as the student progresses through three iterative training stages. This provides strong empirical evidence that our iterative weak-to-strong strategy enhances model capacity through progressive self-teaching. Notably, substantial gains are achieved on challenging benchmarks where existing models struggle: after three iterations, relative SRCC improvements of $30.59\%$, $20.55\%$, and $8.27\%$ are obtained on LIVE-YT-HFR, Waterloo-IVC-4K, and KVQ, respectively.
Comparison with SOTAs. We compare our Stage 3 student model with state-of-the-art baselines. Our model surpasses all competitors, including the five teacher models and two recent LMM-based approaches, VQA2 (Jia et al., 2024) and VQAThinker (Cao et al., 2025). Notably, VQA2 is trained on over 157k labeled samples, while VQAThinker leverages reinforcement learning with advanced LMM backbones. In contrast, our weak-to-strong learning strategy achieves state-of-the-art performance without any human-labeled data, underscoring its effectiveness and practical value.
5 Discussion
Developing generalized VQA models remains a fundamental challenge due to the vast diversity of real-world distortions and the strong influence of video content. Supervised learning on human-labeled data cannot feasibly cover this space, highlighting the urgent need for unsupervised and weakly supervised paradigms. In this work, we demonstrate that it is possible to learn from weak VQA models and even surpass their performance. Building on this insight, we propose a framework that integrates diverse homogeneous and heterogeneous VQA teachers through a learning-to-rank formulation, and further enhances generalization via an iterative W2S training strategy, where progressively stronger students are recycled as new teachers. This design enables cumulative transfer of knowledge beyond any single teacher and drives the model’s self-evolution toward increasingly generalized quality assessment.
Looking forward, this paradigm suggests a pathway toward scalable VQA foundation models. The community can leverage a broad spectrum of supervision sources, leveraging expert-domain VQA models (e.g., VMAF for video compression), utilizing powerful LMMs with carefully designed prompt engineering, and employing text-to-video generation algorithms to synthesize videos of varying quality through specified prompts. By unifying these heterogeneous signals, future research may move toward constructing foundation models for VQA that generalize across content domains, distortion types, and application scenarios—ultimately serving as universal quality assessors for both natural and generative videos.
@@ 6 Conclusion
This paper introduces a weak-to-strong (W2S) paradigm for video quality assessment that leverages multiple weak teachers and iterative self-teaching to train stronger students without relying on human annotations. Through the integration of homogeneous and heterogeneous teachers under a ranking-based formulation, and the use of iterative W2S training, our approach consistently surpasses the teacher models across ten benchmarks, with particularly strong gains on challenging out-of-distribution benchmarks. The results highlight the potential of W2S as a scalable and effective alternative to traditional annotation-dependent training pipelines.
LLM Usage Statement. Large language models are used to aid in polishing the writing of this paper, but they are not involved in the research design, experimental process, or analysis.
Ethics Statement. All videos used in this work are obtained through a filtering pipeline that ensures only publicly available content with permissive licenses is included.
Reproducibility Statement. Detailed descriptions of the data processing pipeline, training and inference configurations are provided in the main paper and appendix. Our code link: https://github.com/clh124/W2S-VQA.
Appendix A More Details of Our $D_{\text{w2s}}$ Database
@@ A.1 Analysis of the Collected Videos
As shown in Fig. 8, our dataset is collected from multiple popular social media platforms with relatively uniform sampling, comprising $20\%$ from Bilibili, $20\%$ from Youku, $25\%$ from YouTube, and $35\%$ from TikTok. All videos are obtained through a filtering pipeline that ensures only publicly available content with permissive licenses is included. Notably, our dataset covers a diverse range of content categories, exceeding twenty in total. In addition to common categories such as lifestyle, food, and animals, it also includes specialized categories such as gaming, AI-generated content, and high-resolution content. To illustrate the diversity of our dataset, we present a variety of video samples in Fig. 7, showcasing the broad range of content available in our large-scale video quality assessment (VQA) dataset. Unlike existing datasets, which often focus on specific formats, our dataset encompasses a wider variety of formats, including both landscape and portrait orientations, as well as various resolutions. This diversity enhances the comprehensiveness of our dataset, making it more suitable for evaluating video quality across a wide kinds of scenarios. A detailed breakdown of our database, including pair types and the corresponding number of videos, is provided in Table 2.
@@
Table 2: Statistics of raw videos and video pairs in the $D_{\text{w2s}}$ dataset.
Category Subtype Videos Video Pairs
$D_{\text{w2s}}^{(1)}$ $D_{\text{w2s}}^{(2)}$ $D_{\text{w2s}}^{(3)}$ $D_{\text{w2s}}^{(1)}$ $D_{\text{w2s}}^{(2)}$ $D_{\text{w2s}}^{(3)}$
Ensembling homogeneous teachers - 200k 100k 50k 250k 85k 85k
Integrating heterogeneous teachers Spatial 50k 2k 2k 160k 5k 5k
Temporal 20k 1k 1k 40k 5k 5k
Compression 10k 1k 1k 50k 5k 5k
Total 280k 384k 438k 500k 600k 700k
@@ A.2 Analysis of Low-level Metrics
Our data selection strategy is based on a mixed-integer programming method (Vonikakis et al., 2017), which optimizes dataset composition by aligning feature histograms. Specifically, we utilize this approach to match the distributions of nine low-level metrics (blockiness (Romaniak et al., 2012), blur (Narvekar & Karam, 2011), contrast (Peli, 1990), noise, flickering (Pandel, 2008), colourfulness (Hasler & Suesstrunk, 2003), luminance, spatial information (SI)  (ITU-T P.910, 2008), and temporal information (TI) (ITU-T P.910, 2008)) between our dataset and the LSVQ dataset. Each metric is computed as follows:
Figure 9: Distribution of nine metrics on the LSVQ dataset, as well as on our dataset before and after sampling.
@@
(Romaniak et al., 2012) is quantified by analyzing the luminance differences between pixels within and across encoding blocks. Specifically, we compute the absolute luminance differences between adjacent pixel pairs within the same encoding block (internal pixel pairs) and those spanning adjacent blocks (external pixel pairs). The blockiness metric is then determined as the ratio of the total sum of internal pixel difference values to the total sum of external pixel difference values across the entire video frame:
@@
is measured using the Cumulative Probability of Blur Detection (CPBD) (Narvekar & Karam, 2011), which evaluates perceptual sharpness based on edge width distribution. A higher CPBD value indicates a sharper image. Given an edge pixel $e_{i}$, its width $w(e_{i})$ is compared with the Just Noticeale Blur (JNB) threshold, determining the blur detection probability $w_{JNB}(e_{i})$. The final CPBD score is computed as:
@@
is a measure of the dispersion of pixel intensity values within the video frame and can be quantified using the standard deviation of grayscale intensities (Peli, 1990). Specifically, for a grayscale image $I(x,y)$, the mean intensity $\mu$ is first computed as:
@@
occurs when an encoder skips macroblocks to conserve bitrate, especially in low-texture, slow-motion regions (Pandel, 2008). It is quantified by counting macroblock transitions from an “unupdated” to an “updated” state, with a threshold $T_{f}$ ensuring only significant changes are considered. The flickering metric is computed as:
@@
quantifies color distribution differences across RGB channels, following (Hasler & Suesstrunk, 2003). Given a frame with RGB channels $R,G,B$, we compute:
@@
A.3 More Details on Synthetic Distortion Data
@@ A.3.1 Spatial Distortions
Table 3: An overview of our testing datasets.
KoNViD-1k (Hosu et al., 2017) 2017 1,200 1,200 540p 8 24, 25, 30 In-the-wild
LIVE-VQC (Sinno & Bovik, 2018) 2018 585 585 240p–1080p 10 30 In-the-wild
YouTube-UGC (Wang et al., 2019) 2019 1,380 1,380 360p–4K 20 30 In-the-wild
LSVQ (Ying et al., 2021) 2021 38,811 38,811 99p–4K 5–12 << 60 In-the-wild
Waterloo-IVC-4K (Li et al., 2019b) 2019 1200 20 540p, 1080p, 4k 9-10 24, 25, 30 H.264 compression
LIVE-YT-HFR (Madhusudana et al., 2021) 2021 480 16 1080p 6-10 24, 30, 60, 82, 98, 120 Frame rate, VP9 compression
LIVE-YT-Gaming (Yu et al., 2022) 2022 600 600 360p–1080p 8–9 30, 60 PGC, UGC
CGVDS (Saha et al., 2023) 2023 360 15 480p, 720p, 1080p 30 20, 30, 60 H.264 compression
KVQ (Lu et al., 2024) 2024 4200 600 - 3-8 - UGC
A.3.2 Temporal Distortions
@@ A.4 More Details on Testing Datasets
Appendix B More Details of Quality Annotation
B.1 Weak Models for Pseudo-labeling
Table 4: Comparison of model parameters and architecture.
Model Parameters (M) Architecture
MinimalisticVQA(VII) 86.93 Swin-B
MinimalisticVQA (IX) 121.59 Swin-B + SlowFast
FAST-VQA 29.97 Swin-Tiny
DOVER 58.06 Swin-Tiny + Conv-Tiny
Q-Align 8204.56 mPLUG-Owl2
Our strong model 8075.24 LLaVA-OneVision-Chat + SlowFast
We choose five SOTA VQA models: MinimalisticVQA (VII) (Sun et al., 2024), MinimalisticVQA (IX) (Sun et al., 2024), FAST-VQA (Wu et al., 2022), DOVER (Wu et al., 2023a), and Q-Align (Wu et al., 2023b) as weak teachers to formulate our pseudo quality annotation. The detail introduction of the five models is as follows:
MinimalisticVQA (VII)
employs Swin Transformer-B (Liu et al., 2022), pre-trained on ImageNet-1K (Deng et al., 2009), as the spatial quality analyzer to extract quality-aware spatial features from key frames, ensuring robust spatial quality assessment.
MinimalisticVQA (IX)
builds upon MinimalisticVQA (VII) by incorporating a temporal quality analyzer to account for motion distortions. The temporal quality analyzer, implemented using the SlowFast (Feichtenhofer et al., 2019) network pre-trained on the Kinetics-400 (Carreira & Zisserman, 2017) dataset, extracts motion-related features from video chunks, enhancing the model’s ability to assess temporal quality variations.
@@ B.2 Prompts for Model Training
We construct the label prompts for our large-scale dataset using a fixed template. For the single-video input:
Question: "You will now receive a video: <image>. Please watch the video carefully and answer the following question: What is your overall rating of the quality of this video?"
Answer: "[quality score]"
For the dual-video input:
Question: "You will now receive two videos. The first video: <image>. The second video: <image>. Please watch both videos carefully and answer the following question: Compared to the first video, how would you rate the quality of the second video?"
Answer: "The quality of the second video is [level] compared to the first video."
Appendix C More Details of Our Strong student Model
C.1 Model Structure
As illustrated in Fig. 3, our model comprises three components: a visual feature extractor, a text tokenizer, and an LLM decoder.
$\bm{r}$ $=\mathcal{L}(\bm{f}^{v}_{A},\bm{f}^{v}_{B},\bm{f}^{p}),$ (22)
where $\bm{r}$ is expected to belong to {“superior”, “better”, “similar”, “worse”, “inferior”}.
C.2 Training Details
C.2.1 Training Setup
The model is trained using the DeepSpeed framework with mixed-precision floating-point operations to optimize memory and computational efficiency. The training is conducted for one epoch with a batch size of 2 per device and a gradient accumulation step of 1. The optimizer follows AdamW with a initial learning rate of $1\times 10^{-4}$, a cosine learning rate schedule, and a warm-up ratio of 0.03.
@@ C.2.1 Training Setup
C.2.2 Auxiliary Confidence Loss
As mentioned in Section 4.3, we introduce an auxiliary confidence loss to encourage the model to maintain high-confidence predictions, especially in the presence of noisy weak supervision. The final training objective is a dynamically weighted combination of the cross-entropy loss $\mathcal{L}_{\text{CE}}$ and the confidence loss $\mathcal{L}_{\text{conf}}$:
$\mathcal{L}=(1-\lambda)\cdot\mathcal{L}_{\text{CE}}+\lambda\cdot\mathcal{L}_{\text{conf}},$ (23)
where $\lambda$ is an adaptive weighting factor that balances between trusting the weak labels and relying on the model’s own confidence. The confidence loss is defined as the average entropy over the predicted token probability distributions:
$\mathcal{L}_{\text{conf}}=\frac{1}{N}\sum_{i=1}^{N}H(p_{\theta}(x_{i}))=-\frac{1}{N}\sum_{i=1}^{N}\sum_{c}p_{\theta}(c|x_{i})\log p_{\theta}(c|x_{i}),$ (24)
where $p_{\theta}(c|x_{i})$ denotes the predicted probability of vocabulary token $c$ given input $x_{i}$. By minimizing the entropy of the predicted distribution, we encourage the model to produce more confident next-token predictions.
To dynamically adjust $\lambda$ during training, we introduce a temperature-based confidence estimation mechanism. Specifically, we define:
$\lambda=\alpha\cdot\min\left(1.0,\frac{t}{T_{\text{warmup}}}\right),$ (25)
where $t$ denotes the current training step ratio (normalized to $[0,1]$), and $T_{\text{warmup}}$ is the warm-up period, which we set to $10\%$ of the total training steps. This warm-up phase ensures that the strong model gradually learns to rely on its own confidence, while initially being guided by the weak labels. The factor $\alpha$ is computed as the ratio between the temperature-scaled exponentials of the two losses:
$\alpha=\frac{\exp(\mathcal{L}_{\text{conf}}/T)}{\exp(\mathcal{L}_{\text{conf}}/T)+\exp(\mathcal{L}_{\text{CE}}/T)}.$ (26)
Here, $T$ is a temperature parameter that controls the sharpness of the weighting between the two loss components. We linearly decrease $T$ from $0.5$ to $0.1$ during the warm-up period to gradually increase the sensitivity of $\alpha$ to differences in the two loss values.
C.3 Inferring Details
C.3.1 Probability Modeling

### v4

@@ 5.2 Performance Analysis
Iterative Self-Improvement Training Strategy. Consistent performance improvements are observed across both in-domain and OOD datasets as the model advances through training stages. This empirical evidence demonstrates that our iterative training strategy effectively enhances model capability through progressive self-learning. Notably, substantial gains are achieved on challenging VQA benchmarks where existing models struggle: the LIVE-YT-HFR, Waterloo-IVC-4K, and KVQ datasets exhibit relative improvements of $16.2\%$, $21.4\%$, $9.3\%$ in SRCC respectively, after three-stage refinement.
In-Domain Performance Comparison. We observe that all proposed model variants outperform the five competing VQA models, which are used for label annotation in Section 3.2.1, confirming our self-supervised approach successfully distills ensemble knowledge from its teacher models on unlabeled data. However, our models are slightly outperformed by FAST-VQA and DOVER on LIVE-VQC, which exhibits more complex temporal distortions compared to other datasets [52]. We attribute this performance gap to architectural differences: FAST-VQA and DOVER fine-tune a 3D DNN (i.e., VideoSwin) specifically optimized for spatiotemporal feature extraction, whereas our framework handles temporal information using a pre-trained SlowFast model.
OOD Performance Comparison. Our models significantly outperform competing approaches on the OOD benchmarks, achieving a $24.7\%$ improvement in SRCC for overall OOD evaluation compared to the competing approaches. We attribute the strong OOD performance to two key factors. First, our training datasets include compression distortions (H.264 and H.265), and the synthetic quality ranking labels provide useful supervision for assessing compression-related video quality for these datasets. Second, our iterative self-improvement training strategy enables progressive adaptation to unseen distortion types, such as frame rate inconsistencies in LIVE-YT-HFR, AVS/VP9 compression artifacts in Waterloo-IVC-4K, and enhancement-induced distortions in KVQ—none of which are present in the training data. These results demonstrate the effectiveness of the proposed self-supervised VQA framework in real-world quality assessment scenarios.
5.3 Supervised Performance Comparison
We conduct supervised fine-tuning of our self-supervised trained model on the in-domain VQA benchmarks, as shown in Table 3. The results reveal that: (1) Our fine-tuned model surpasses all competing VQA baselines with 3.5% higher SRCC, and (2) it achieves $6.4\%$ improvement over the baseline trained without self-supervised initialization. These results confirm that self-supervised representation learning critically enhances downstream fine-tuning effectiveness.
Furthermore, we compare the prediction accuracy of pairwise comparisons in LMM-PVQA with four open-source LMMs that have demonstrated strong performance in video understanding. As shown in Table 6, LMM-PVQA significantly outperforms these advanced LMMs, providing highly accurate quantitative predictions. In contrast, other models primarily rely on high-level instruction-tuning datasets, resulting in suboptimal accuracy in quality prediction. This suggests that existing LMMs exhibit limited quality perception capabilities, highlighting the advantages of our proposed approach in video quality assessment.
5.4 Ablation Study
We conduct the ablation experiments to validate the effectiveness of LMM-PVQA, with results shown in Table 4.
Motion Encoder. Incorporating the motion encoder for temporal distortion representation yields consistent improvements across all benchmarks, particularly achieving $33.5\%$ SRCC improvement on LIVE-YT-HFR—the benchmark specifically designed for high frame-rate distortion analysis. This verifies that explicit motion modeling is essential for capturing temporal-related distortions.
Synthetic Distortion Data. The incorporation of the PLVD-DS dataset also demonstrates consistent performance improvements across all benchmarks, yielding marginal gains on in-domain benchmarks while achieving substantial enhancements on OOD benchmarks. It can be attributed that the PLVD-VQA dataset have already equipped the model with quality assessment capabilities for in-domain content, whereas the PLVD-DS dataset help mitigate the domain gap on OOD benchmarks by simulating common artificial degradations (e.g., compression).
Iterative Self-Improvement Training Strategy. Since our iterative self-improvement training strategy introduces new training samples, it is critical to verify that performance gains stemming from the strategy itself rather than additional data. To isolate this effect, we train a Stage 2 model using the same base training set (PLVD-Part1/Part2), where videos in PLVD-Part2 are labeled by five VQA judges, without incorporating the Stage 1 model as the new judge. Experimental results confirm that merely increasing training data (without iterative refinement) fails to surpass the Stage 1 model’s performance, thereby validating the effectiveness of our strategy in enabling self-improvement through iterative optimization.
@@ 6 Conclusion
We propose a self-supervised learning framework for VQA to mitigate dependence on human-annotated datasets. Our approach constructs a large-scale video pair dataset labeled via established VQA models and synthetic distortion simulations, then adopts a learning-to-rank paradigm to learn quality assessment capabilities. By integrating an iterative self-improvement training strategy, our model progressively enhances its evaluation performance through self-learning. Our method achieves state-of-the-art zero-shot performance on both in-domain and OOD benchmarks, demonstrating its effectiveness.
Limitations. Our method is inherently designed for pairwise quality comparison. To obtain absolute quality scores for a single video, it must be compared against multiple anchor videos (five in this study), which significantly increases inference time.
Future Work. We plan to explore more automated video pair annotation strategies, such as leveraging expert-domain VQA models (e.g., VMAF for video compression), utilizing LMMs with carefully designed prompt engineering, and employing text-to-video generation algorithms to synthesize videos of varying quality through specified prompts, thus substantially expanding the diversity and scale of training data. Furthermore, we intend to extend our framework to incorporate additional modalities, such as images and audio, to develop a more generalizable quality assessment model.
Supplementary Material
Figure 4: Examples videos of different categories in our large dataset.
Appendix A More Details of Our PLVD Database
@@ A.1 Analysis of the Collected Videos
As shown in Fig. 5, our dataset is collected from multiple popular social media platforms with relatively uniform sampling, comprising $20\%$ from Bilibili, $20\%$ from Youku, $25\%$ from YouTube, and $35\%$ from TikTok. Notably, our dataset covers a diverse range of content categories, exceeding twenty in total. In addition to common categories such as lifestyle, food, and animals, it also includes specialized categories such as gaming, AI-generated content, and high-resolution content. To illustrate the diversity of our dataset, we present a variety of video samples in Fig. 4, showcasing the broad range of content available in our large-scale video quality assessment (VQA) dataset. Unlike existing datasets, which often focus on specific formats, our dataset encompasses a wider variety of formats, including both landscape and portrait orientations, as well as various resolutions. This diversity enhances the comprehensiveness of our dataset, making it more suitable for evaluating video quality across a wide kinds of scenarios.
@@ A.2 Analysis of Low-level Metrics
Our data selection strategy is based on a mixed-integer programming method [56], which optimizes dataset composition by aligning feature histograms. Specifically, we utilize this approach to match the distributions of nine low-level metrics (blockiness [45], blur [41], contrast [43], noise, flickering [42], colourfulness [14], luminance, spatial information (SI) [19], and temporal information (TI) [19]) between our dataset and widely used VQA datasets. Each metric is computed as follows:
@@
[45] is quantified by analyzing the luminance differences between pixels within and across encoding blocks. Specifically, we compute the absolute luminance differences between adjacent pixel pairs within the same encoding block (internal pixel pairs) and those spanning adjacent blocks (external pixel pairs). The blockiness metric is then determined as the ratio of the total sum of internal pixel difference values to the total sum of external pixel difference values across the entire video frame:
@@
is measured using the Cumulative Probability of Blur Detection (CPBD) [41], which evaluates perceptual sharpness based on edge width distribution. A higher CPBD value indicates a sharper image. Given an edge pixel $e_{i}$, its width $w(e_{i})$ is compared with the Just Noticeale Blur (JNB) threshold, determining the blur detection probability $w_{JNB}(e_{i})$. The final CPBD score is computed as:
@@
is a measure of the dispersion of pixel intensity values within the video frame and can be quantified using the standard deviation of grayscale intensities [43]. Specifically, for a grayscale image $I(x,y)$, the mean intensity $\mu$ is first computed as:
@@
occurs when an encoder skips macroblocks to conserve bitrate, especially in low-texture, slow-motion regions [42]. It is quantified by counting macroblock transitions from an “unupdated” to an “updated” state, with a threshold $T_{f}$ ensuring only significant changes are considered. The flickering metric is computed as:
@@
quantifies color distribution differences across RGB channels, following [14]. Given a frame with RGB channels $R,G,B$, we compute:
@@
A.3 More Details of PLVD-DS Datas
@@ A.3.1 Spatial Distortion
Table 5: An overview of our testing datasets
KoNViD-1k [16] 2017 1,200 1,200 540p 8 24, 25, 30 In-the-wild
LIVE-VQC [50] 2018 585 585 240p–1080p 10 30 In-the-wild
YouTube-UGC [60] 2019 1,380 1,380 360p–4K 20 30 In-the-wild
LSVQ [70] 2021 38,811 38,811 99p–4K 5–12 << 60 In-the-wild
Waterloo-IVC-4K [26] 2019 1200 20 540p, 1080p, 4k 9-10 24, 25, 30 H.264 compression
LIVE-YT-HFR [33] 2021 480 16 1080p 6-10 24, 30, 60, 82, 98, 120 Frame rate, VP9 compression
LIVE-YT-Gaming [71] 2022 600 600 360p–1080p 8–9 30, 60 PGC, UGC
CGVDS [47] 2023 360 15 480p, 720p, 1080p 30 20, 30, 60 H.264 compression
KVQ [32] 2024 4200 600 - 3-8 - UGC
@@
A.3.2 Temporal Distortion
@@ A.4 More Details on Testing Datasets
Table 6: Performance comparison of our model in terms of prediction accuracy on four VQA datasets against advanced LLMs. Additionally, performance comparison of an ensemble of five methods versus a single model on the same datasets
Method $_{\textbf{test}} LSVQ$$_{\textbf{1080p}}$ KoNViD-1k LIVE-VQC YouTube-UGC
LLaVA-NeXT-Video (8B) [74] 0.508 0.531 0.512 0.531 0.506
mPLUG-Owl3-8B [68] 0.547 0.534 0.605 0.548 0.543
InternVL2_\_5-8B [59] 0.566 0.629 0.583 0.596 0.588
Qwen2.5-7B-VL [2] 0.696 0.666 0.740 0.664 0.631
LLaVA-ov-chat [24] 0.667 0.641 0.727 0.664 0.639
MinimalisticVQA(VII) [52] 0.907 0.817 0.906 0.839 0.784
MinimalisticVQA(IX)  [52] 0.926 0.845 0.904 0.843 0.820
FAST-VQA [61] 0.918 0.836 0.918 0.832 0.765
DOVER [64] 0.922 0.840 0.915 0.860 0.812
Q-Align [65] 0.922 0.849 0.923 0.846 0.822
Ensemble five methods 0.939 0.864 0.935 0.860 0.831
Soft Ranking (Stage 1) 0.926 0.855 0.935 0.849 0.853
Appendix B More Details of Pairwise Quality Annotation
B.1 VQA Models for Pseudo-labeling
We choose five SOTA VQA models: Minimalistic-VQA (VII) [52], Minimalistic-VQA (IX) [52], FAST-VQA [61], DOVER [64], and Q-Align [65] as our initial judges to formulate our pairwise quality annotation. As shown in Table 6, the ranking accuracy of the five ensemble methods is higher than that of any single model. The detail introduction of the five models is as follows:
Minimalistic-VQA (VII)
employs Swin Transformer-B [31], pre-trained on ImageNet-1K [9], as the spatial quality analyzer to extract quality-aware spatial features from key frames, ensuring robust spatial quality assessment.
Minimalistic-VQA (IX)
builds upon Minimalistic-VQA (VII) by incorporating a temporal quality analyzer to account for motion distortions. The temporal quality analyzer, implemented using the SlowFast [12] network pre-trained on the Kinetics-400 [5] dataset, extracts motion-related features from video chunks, enhancing the model’s ability to assess temporal quality variations.
@@ B.2 Prompts for Model Training
We final construct the label prompt of our large dataset using a fixed template:
Question: "Now you will receive two videos. The first video:<image>. The second video:<image>. Please watch these videos carefully, and then answer the following question: Comparing with the first video, how do you assess the quality of the second video?"
Answer: "The quality of the second video is [level] to the first video."
Appendix C More Details of Our Model
C.1 Training Details
The model is trained using the DeepSpeed framework with mixed-precision floating-point operations to optimize memory and computational efficiency. The training is conducted for one epoch with a batch size of 1 per device and a gradient accumulation step of 1. The optimizer follows AdamW with a learning rate of $1\times 10^{-5}$, a cosine learning rate schedule, and a warmup ratio of 0.03.
@@ C.1 Training Details
C.2 Inferring Details
C.2.1 Probability Modeling